OpenStudy (anonymous):

Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (ja1):

\[4x - 7 = 9\]

OpenStudy (ja1):

now lets solve it together so we can find out if its okay or not

OpenStudy (ja1):

try and follow the steps described in the question.

OpenStudy (ja1):

\[4x−7=9\]

OpenStudy (ja1):

@marrero

OpenStudy (anonymous):

h ja1

OpenStudy (anonymous):

@JA1

OpenStudy (ja1):

hey, so do you know how to solve it?

OpenStudy (anonymous):

ight lets see

OpenStudy (anonymous):

7 + 7 = 14 & 7 + 9 = 16

OpenStudy (ja1):

okay good so the steps would be formatted like this: 4x - 7 = 9 +7 +7 4x = 16 (because -7 and +7 cancel out) now you divide both sides by 4 like it said above: 4x = 16 /4 /4 x = 4 (4*4 = 16) so your answer is x = 4 but you can change 9 into any number as long as the format is 4x - 7 = #

OpenStudy (ja1):

do you understand why all this is true?

OpenStudy (anonymous):

nope not realy

OpenStudy (ja1):

Ok so the question asked for an equation that you have to add 7 and divide by 4, and to solve the equation i made (4x - 7 =9) you have to first cancel out the 7 so you add 7 to both sides, and you have to isolate the variable by dividing the 4 next to it.

OpenStudy (ja1):

so it fits the description, the = 9 can be anything it can be =10 or =13 or even =100 but the equation must have 4x - 7 first

OpenStudy (anonymous):

oh,ok

OpenStudy (anonymous):

thx JA1 ur my hero!

OpenStudy (ja1):

yah if you have any more questions feel free to ask me, i can explain further (and better) xD

OpenStudy (anonymous):

ight